A firm uses 100 hours of labor and 20 units of capital to produce 5,000 bicycles a day. Labor’s marginal product is 4 bicycles per day and the marginal product of capital is 10 bicycles per day. Each unit of labor costs $20 and each unit of capital costs $40 per hour. If the firm wants to continue producing 5,000 bicycles per day at the lowest possible cost, it should

A. Continue using 100 hours of labor and 20 units of capital.

B. Use less of both inputs

C. Use more capital and less labor

D. Use more labor and less capit
Answer #1

Ans.- (C)

Explanation-

At the optimal combination of inputs we have :

RTS = w/ r ( where RTS = MPl/MPk and w = price of labor and r = price of capital)

So, RTS = 4/10 = 0.4

w/r = 20/40 = 0.5

Since RTS = MPl/MPk<w/r

So, MPl/w <MPk/r

So, firm should use more capital and less labor.